# An upper bound of the silhouette validation metric for clustering

### v2

@@ 4 A data-dependent upper bound of the silhouette
We propose a data-dependent ceiling on the ASW that can be computed before any clustering is performed. For each data point, we derive a sharp upper bound of its silhouette width. Here, sharp means that the individual silhouette width of a data point can never exceed the bound, and there exists a clustering in which the bound is exactly achieved. Averaging these per-point ceilings produces a single, dataset-level upper bound on the ASW. The bound is computed from the dissimilarity matrix of the dataset in $O(n^{2}\log n)$ time, and the proposed algorithm supports optional minimum cluster size constraints to further tighten the bound.
Consider an arbitrary clustering $\mathcal{C}_{K}$ of the dataset $\{1,\ldots,n\}$ with dissimilarites $\Delta$. We will now formalize the upper bound of the silhouette width for point $i$, i.e. $s(i|\mathcal{C}_{K},\Delta)$.
For data point $i$, let $|C_{I}|$ denote the size of its cluster in $\mathcal{C}_{K}$. The average distance between $i$ and the remaining data points in the same cluster is not less than the average distance between $i$ and the $|C_{I}|-1$ data points closest to $i$. Similarly, the average distance between $i$ and the $n-|C_{I}|$ points farthest from $i$ is not smaller than the average distance between $i$ and the data points in the neighboring cluster closest to $i$. These two observations are combined to construct a value that is guaranteed to be greater than or equal to $s(i|\mathcal{C}_{K},\Delta)$. We formalize this reasoning through the $\Lambda$-quotient of point $i$.
@@
Let $\Lambda\in\mathbb{N}$ with $1\leq\Lambda\leq n-1$. The $\Lambda$-quotient of data point $i$ is
$q(i,\Delta,\Lambda)\coloneqq\begin{cases}\dfrac{n-\Lambda}{\Lambda-1}\cdot\dfrac{\sum_{j=1}^{\Lambda-1}\hat{\Delta}_{ij}}{\sum_{j=\Lambda}^{n-1}\hat{\Delta}_{ij}}&\textnormal{if }\Lambda>1,\\[4.30554pt] 1,&\textnormal{if }\Lambda=1.\end{cases}$
@@
$f(i,\Delta)\coloneqq\min_{1\leq\Lambda\leq n-1}\{q(i,\Delta,\Lambda)\}$
denote a minimal $\Lambda$-quotient. Then
@@
$ASW(\mathcal{C}^{*},\Delta)\leq 1-\frac{1}{n}\sum_{i}f(i,\Delta).$ (1)
We remark that the right-hand side in 1 depends only on the underlying data $\Delta$. Crucially, we note that $1-\frac{1}{n}\sum_{i}f(i,\Delta)\leq 1$. Moreover, since any $\Lambda$-quotient naturally corresponds to a $2$-clustering of the dataset, the bound in Proposition 1 is indeed sharp.
Finally, imposing the narrower definition
$f_{\kappa}(i,\Delta)\coloneqq\min_{\kappa\leq\Lambda\leq n-\kappa}\{q(i,\Delta,\Lambda)\}$
for some $\kappa\in\{1,\ldots,\lfloor n/2\rfloor\}$ is equivalent to constraining the solution space to clusterings having $\min_{I}|C_{I}|\geq\kappa$. Often, as we shall see, this leads to strictly tighter limits.
@@ 4.1 Algorithm
Input: $\kappa$;
// Smallest allowed cluster size, $1\leq\kappa\leq\lfloor n/2\rfloor$
@@
Output: $A\in\mathbb{R}^{n}$;
// Array of individual upper bounds
@@ 8 $x\leftarrow 0$ ;
 // Initialize $\Lambda$-quotient
9 if $\kappa>1$ then
10    $x\leftarrow\sum_{j=1}^{\kappa-1}\hat{\Delta}_{i,j}$ ;
11    $q\leftarrow\frac{x}{\kappa-1}\big/\frac{y}{n-\kappa}$;
@@
13 for $\Lambda=\kappa+1,\ldots,n-\kappa$ do
14    $x\leftarrow x+\hat{\Delta}_{i,\Lambda-1}$;
15    $y\leftarrow y-\hat{\Delta}_{i,\Lambda-1}$;
16    $q_{\text{candidate}}\leftarrow\dfrac{x}{\Lambda-1}\big/\dfrac{y}{n-\Lambda}$;
@@
       // Update minimal $\Lambda$-quotient
@@ 20  Append 1q1-q to AA;
The workflow for computing the proposed upper bound is illustrated in Fig.4. Following this workflow, Algo. 1 computes for each data point $i$ the quantity $1-f_{\kappa}(i,\Delta)$, which is a sharp upper bound of the silhouette width of $i$ in the solution space $\min_{I}|C_{I}|\geq\kappa$. The array $A$ collects these values across all points, and its average equals the dataset-level upper bound.
Time complexity
The formation of $\hat{\Delta}$ requires sorting each of the $n$ rows, which costs $\mathcal{O}(n^{2}\log n)$. The subsequent linear scans are $\mathcal{O}(n^{2})$, so the overall run time is $\mathcal{O}(n^{2}\log n)$. This is slightly slower than computing the $\mathrm{ASW}$, which takes $\mathcal{O}(n^{2})$ time [medoid].
@@
Given $\Delta$ and $\kappa=1$, let $\texttt{UB}(\Delta)$ denote the mean of the output array $A$ from Algo. 1. Then $\texttt{UB}(\Delta)$ corresponds to the right-hand side of Eq. (1). We refer to this as the global upper bound. Choosing $\kappa>1$ imposes the restriction that no cluster is allowed to contain fewer than $\kappa$ objects. Let $\texttt{UB}_{\kappa}(\Delta)$ denote the $\mathrm{ASW}$ upper bound for this constrained solution space. We will refer to $\texttt{UB}_{\kappa}(\Delta)$ as an adjusted upper bound.
Remark
The upper bound $\texttt{UB}(\Delta)$ is not guaranteed to be close to the true global maximum of $\mathrm{ASW}$. Its tightness is data-dependent.
Use cases
In theory, when sharp, the upper bound $\texttt{UB}(\Delta)$ can be used to confirm that an empirical ASW has reached the global optimum. Even when loose, it increases the interpretability of empirical measurements.
In real-world scenarios where the global upper bound appears loose, we speculate that adjusted versions can provide more practical use.
The individual silhouette width upper bounds generated by Algo. 1 provide a neat complement to standard silhouette visualizations, where the point-wise silhouette widths are typically displayed next to each other in a bar graph.
In the Experiments section, we examine these use cases.
Figure 4: Workflow for computing the proposed $\mathrm{ASW}$ upper bound.
@@ 4.2 Extension to the macro-averaged silhouette
The standard $\mathrm{ASW}$ can produce misleading results when the clusters vary greatly in size. In [john], John Pavlopoulos, Georgios Vardakas, and Aristidis Likas show that the so-called macro-averaged silhouette is more robust to cluster-size imbalance. For clustering $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$, the macro-averaged silhouette is defined as
@@
Consider an empirical clustering result $\widehat{\mathcal{C}}_{K}=\{\widehat{C}_{I}\}_{I=1}^{K}$. We narrow our scope in this context to include all clusterings whose cluster sizes match those of $\widehat{\mathcal{C}}_{K}$. This narrower solution space is still generally astronomically large. Thus, an efficiently computable data-dependent upper bound is of interest as a relevant benchmark. Let $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ be any clustering from our constrained solution space, and let $\texttt{ub}_{i}$ denote the silhouette width upper bound of data point $i$, i.e. $s(i|\mathcal{C}_{K},\Delta)\leq\texttt{ub}_{i}$. Moreover, let $\{\texttt{ub}_{\sigma(i)}\}_{i=1}^{n}$ denote the upper bounds sorted in ascending order, and let $\{N_{I}\}_{I=1}^{K}$ denote the assumed cluster sizes in descending order. With $N_{0}\coloneqq 1$, it follows from the rearrangement inequality that
@@
The last expression is independent of any particular cluster assignment and therefore provides a valid upper bound on the macro-averaged silhouette across all clusterings within the constrained solution space. This formulation is particularly useful in applications where cluster sizes are known or can be reasonably estimated. Even in the absence of such assumptions, the resulting bound serves as a sharper and more informative benchmark for evaluating empirical clustering results.
5 Experiments, results & discussion

### v4

@@ 4 A data-dependent upper bound of the silhouette
We propose a data-dependent ceiling on the $\operatorname{ASW}$ that can be computed before any clustering is performed. For each data point, we derive an upper bound of its silhouette width. Averaging these per-point ceilings produces a single, dataset-level upper bound of the $\operatorname{ASW}$. The bound is computed from the dissimilarity matrix of the dataset in $O(n^{2}\log n)$ time, and the proposed algorithm supports optional minimum cluster size constraints to further tighten the bound.
Consider an arbitrary clustering $\mathcal{C}_{K}$ of the dataset $\{1,\ldots,n\}$ with dissimilarites $\Delta$. We will now formalize the upper bound of the silhouette width for point $i$, i.e. $s(i|\mathcal{C}_{K},\Delta)$. For our purposes, it will be convenient to work with the off-diagonal entries of $\Delta$, sorted row-wise in ascending order. We denote this transformed dissimilarity matrix by $\hat{\Delta}$ ($\hat{\Delta}\in\mathbb{R}^{n\times n-1}$). In other words, we have for every $i$ $\{\Delta_{ij}\}_{j=1}^{n}\setminus\{\Delta_{ii}\}=\{\hat{\Delta}_{ij}\}_{j=1}^{n-1}$ and $\hat{\Delta}_{i1}\leq\hat{\Delta}_{i2}\leq\cdots\leq\hat{\Delta}_{i(n-1)}$. Constructing $\hat{\Delta}$ requires sorting each of the $n$ rows of $\Delta$, which costs $\mathcal{O}(n^{2}\log n)$ time in total.
For data point $i$, let $|C_{I}|$ denote the size of its cluster in $\mathcal{C}_{K}$. The average distance between $i$ and the remaining data points in the same cluster is not less than the average distance between $i$ and the $|C_{I}|-1$ data points closest to $i$. Similarly, the average distance between $i$ and the $n-|C_{I}|$ points farthest from $i$ is not smaller than the average distance between $i$ and the data points in the neighboring cluster closest to $i$. These two observations are combined to construct a value that is guaranteed to be greater than or equal to $s(i|\mathcal{C}_{K},\Delta)$. We formalize this reasoning through the $k$-quotient of point $i$.
@@
Let $k\in\mathbb{N}$ with $1\leq k\leq n-1$. The $k$-quotient of data point $i$ is
$q(i,\Delta,k)\coloneqq\begin{cases}\dfrac{n-k}{k-1}\cdot\dfrac{\sum_{j=1}^{k-1}\hat{\Delta}_{ij}}{\sum_{j=k}^{n-1}\hat{\Delta}_{ij}}&\textnormal{if }k>1,\\[4.30554pt] 1,&\textnormal{if }k=1.\end{cases}$
@@
$f(i,\Delta)\coloneqq\min_{1\leq k\leq n-1}\{q(i,\Delta,k)\}$
denote a minimal $k$-quotient. Then
@@
$\operatorname{ASW}(\mathcal{C}^{*},\Delta)\leq 1-\frac{1}{n}\sum_{i}f(i,\Delta).$ (1)
We remark that the right-hand side in 1 depends only on the underlying data $\Delta$. Crucially, we note that $1-\frac{1}{n}\sum_{i}f(i,\Delta)\leq 1$.
We stress that the right-hand side in 1 is not, in general, a sharp upper bound of the ASW. However, for a single data point $i$, the quantity $1-f(i,\Delta)$ is a sharp upper bound of its individual silhouette width in the sense that the individual silhouette width of the data point can never exceed the bound, and there exists a clustering in which the bound is achieved exactly. Indeed, considering the $2$-clustering where $i$ is grouped with its k1k-1 nearest neighbors and all remaining points are grouped into a single separate cluster yields $s(i|C_{K},\Delta)=1-f(i,\Delta)$.
Imposing the narrower definition
$f_{m}(i,\Delta)\coloneqq\min_{m\leq k\leq n-m}\{q(i,\Delta,k)\}$
for some $m\in\{1,\ldots,\lfloor n/2\rfloor\}$ is equivalent to constraining the solution space to clusterings having $\min_{I}|C_{I}|\geq m$. By letting $\mathcal{C}^{*}_{m}$ denote a silhouette-optimal clustering within this solution space, we have $\operatorname{ASW}(\mathcal{C}^{*}_{m},\Delta)\leq 1-\frac{1}{n}\sum_{i}f_{m}(i,\Delta)$.
In real-world applications, analysts sometimes need to enforce minimum cluster size constraints to avoid too small clusters [24, Höppner2008]. From this perspective, the quantity $1-\frac{1}{n}\sum_{i}f_{m}(i,\Delta)$ does not aim to approximate the unconstrained global upper bound in 1, but rather provides an upper bound tailored to the family of clusterings that a practitioner is willing to consider. This allows for a meaningful benchmark in settings where small clusters are inadmissible, and offers flexibility in situations where minimum cluster sizes are part of the problem specification.
@@ 4.1 Algorithm
Input: $m$;
// Smallest allowed cluster size, $1\leq m\leq\lfloor n/2\rfloor$
@@
Output: $S\in\mathbb{R}^{n}$;
// Vector of point-wise upper bounds
@@ 8 $x\leftarrow 0$ ;
 // Initialize $k$-quotient
9 if $m>1$ then
10    $x\leftarrow\sum_{j=1}^{m-1}\hat{\Delta}_{i,j}$;
11    $q\leftarrow\frac{x}{m-1}\big/\frac{y}{n-m}$;
@@
13 for $k=m+1,\ldots,n-m$ do
14    $x\leftarrow x+\hat{\Delta}_{i,k-1}$;
15    $y\leftarrow y-\hat{\Delta}_{i,k-1}$;
16    $q_{\text{candidate}}\leftarrow\dfrac{x}{k-1}\big/\dfrac{y}{n-k}$;
@@
       // Update minimal $k$-quotient
@@ 20  Append 1q1-q to $S$;
Algorithm 1 computes for each data point $i$ the quantity $1-f_{m}(i,\Delta)$, which are sharp upper bounds of the silhouette widths in the solution space $\min_{I}|C_{I}|\geq m$. The vector $S$ collects these values across all points, and its average equals the dataset-level upper bound. As we shall see in the Experiments section, the data point-level upper bounds provide a nice complement to standard silhouette visualizations, where the achieved point-wise silhouette widths are typically displayed next to each other in a bar graph.
Scalability
The formation of $\hat{\Delta}$ requires sorting each of the $n$ rows, which costs $\mathcal{O}(n^{2}\log n)$ time. The subsequent linear scans are $\mathcal{O}(n^{2})$, so the overall runtime is $\mathcal{O}(n^{2}\log n)$. This is slightly slower than computing the $\operatorname{ASW}$, which takes $\mathcal{O}(n^{2})$ time [14]. Since the algorithm operates on the entire $n\times n$ distance matrix, the space complexity of Algorithm 1 is $\mathcal{O}(n^{2})$. This quadratic memory requirement quickly becomes the dominant bottleneck. For example, a dataset with $10^{5}$ samples produces $10^{10}$ pairwise distances; storing these in float32 requires 40 GB of memory. In practice, this limits the algorithm to datasets with at most a few tens of thousands of points on standard hardware. There are ways to mitigate this in large-scale settings. For example, one may approximate the upper bound by computing it on a random subset of the samples. Another possibility is to compute and sort distances on the fly rather than materializing the entire matrix. A more thorough investigation of such scalable variants is an interesting direction for future work, but it lies beyond the scope of this paper.
@@
Given $\Delta$ and $m=1$, let $\texttt{UB}(\Delta)$ denote the mean of the output array $S$ from Algorithm 1. Then $\texttt{UB}(\Delta)$ corresponds to the right-hand side of Eq. (1). We refer to this as the global upper bound. Choosing $m>1$ imposes the restriction that no cluster is allowed to contain fewer than $m$ objects. Let $\texttt{UB}_{m}(\Delta)$ denote the $\operatorname{ASW}$ upper bound for this constrained solution space. We will refer to $\texttt{UB}_{m}(\Delta)$ as a constrained upper bound.
@@ 4.2 Extension to the macro-averaged silhouette
The standard $\operatorname{ASW}$ can produce misleading results when the clusters vary greatly in size. John Pavlopoulos, Georgios Vardakas, and Aristidis Likas [16] show that the so-called macro-averaged silhouette is more robust to cluster size imbalance. For clustering $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$, the macro-averaged silhouette is defined as
@@
Consider an empirical clustering result $\widehat{\mathcal{C}}_{K}=\{\widehat{C}_{I}\}_{I=1}^{K}$. We narrow our scope in this context to include all clusterings whose cluster sizes match those of $\widehat{\mathcal{C}}_{K}$. This narrower solution space is still generally astronomically large. Thus, an efficiently computable data-dependent upper bound is of interest as a relevant benchmark. Let $\mathcal{C}_{K}=\{C_{I}\}_{I=1}^{K}$ be any clustering from our constrained solution space, and let $\texttt{ub}_{i}$ denote the silhouette width upper bound of data point $i$, i.e., $s(i|\mathcal{C}_{K},\Delta)\leq\texttt{ub}_{i}$. Moreover, let $\{\texttt{ub}_{\sigma(i)}\}_{i=1}^{n}$ denote the upper bounds sorted in ascending order, and let $\{N_{I}\}_{I=1}^{K}$ denote the assumed cluster sizes in descending order. With $N_{0}\coloneqq 1$, it follows from the rearrangement inequality that
@@
The last expression is independent of any particular cluster assignment and therefore provides a valid upper bound on the macro-averaged silhouette across all clusterings within the constrained solution space. This formulation is particularly useful in applications where cluster sizes are known or can be reasonably estimated, for example in balanced clustering [15]. Even in the absence of such assumptions, the resulting bound serves as a sharper and more informative benchmark for evaluating empirical clustering results.
Because the macro-averaged silhouette weights clusters equally, it carries the danger of favoring solutions with a few small outlier clusters while masking poor structure in the remaining parts. It should therefore be used only when equal importance of all clusters is appropriate. Moreover, we do not recommend routine comparisons with ASW via heuristics such as $\max(\mathrm{ASW},S_{macro})$. Informed choices should instead be made on a case-by-case basis.
We stress that this extension is not a central focus of the paper. Our aim here is only to illustrate an idea, rather than to conduct a full empirical study of this variant.
5 Experiments
We evaluate the $\operatorname{ASW}$ upper bound by applying it to a variety of datasets, both synthetic and real. Multiple distance metrics are considered to examine the behavior of the upper bound under different notions of dissimilarity. The code used for the experiments is available in the GitHub repository https://github.com/hugo-strang/silhouette-upper-bound which contains all preprocessing steps, dissimilarity constructions, and scripts required to reproduce the reported results.
For a given dissimilarity matrix, an obvious desideratum of the proposed upper bound is that it is relatively close to the maximum $\operatorname{ASW}$. However, the maximum $\operatorname{ASW}$ is of course typically unknown. This presents a fundamental challenge in the evaluation process. To remedy this, we must try to establish a nearly tight lower bound of $\operatorname{ASW}(\mathcal{C}^{*},\Delta)$. To this end, we use PAMSIL clustering by default to obtain an empirical $\operatorname{ASW}$. The closer this value is to the upper bound, the narrower the interval in which the true global maximum lies. PAMSIL is a suitable algorithm since its objective is to maximize $\operatorname{ASW}$. When dealing with larger datasets, we use FastMSC and DynMSC instead.
Given an empirical $\operatorname{ASW}$ value, say $\widehat{S}>0$, the performance metric of interest is the relative error, defined as
$\frac{\operatorname{ASW}(\mathcal{C}^{*},\Delta)-\widehat{S}}{\operatorname{ASW}(\mathcal{C}^{*},\Delta)}=1-\frac{\widehat{S}}{\operatorname{ASW}(\mathcal{C}^{*},\Delta)},$
where $\mathcal{C}^{*}$ denotes a clustering that achieves the global maximum. Since $\operatorname{ASW}(\mathcal{C}^{*},\Delta)$ is unknown in practice, we use the global upper bound $\texttt{UB}(\Delta)$ to obtain a conservative estimate:
$1-\frac{\widehat{S}}{\operatorname{ASW}(\mathcal{C}^{*},\Delta)}~\leq~1-\frac{\widehat{S}}{\texttt{UB}(\Delta)}.$
We refer to this quantity as the worst-case relative error. Its interpretation is straightforward: if the worst-case relative error falls below a tolerance $\epsilon$, then the empirical $\operatorname{ASW}$ is guaranteed to be within $\epsilon$ of the true maximum. Instead of employing this measure to compare clustering algorithms, we will use it to evaluate the tightness of the upper bound relative to the $\operatorname{ASW}$ obtained by the chosen algorithm.
The results of both the global upper bound and the constrained upper bound are reported. In our experiments, we do not impose external constraints on the minimum cluster size. Instead, for each empirical clustering solution (e.g., produced by PAMSIL), we simply record its smallest cluster size and compute $\texttt{UB}_{m}({\Delta})$ using this value of $m$. This approach does not enforce a constraint but allows us to observe how the constrained upper bound behaves under the cluster size patterns that arise in practice.
5.1 Synthetic data
In this section, we look at the proposed upper bound applied to synthetic data generated by the make_blobs function from the scikit-learn library [17]. In this idealized setting, we also demonstrate why a complete scan of the $k$-values is necessary when constructing $f(i,\Delta)$ in Proposition 1. Indeed, it is reasonable to expect that $k=2$ often minimizes $q(i,\Delta,k)$. Lastly, we use synthetic dissimilarity matrices to evaluate the runtime of Algorithm 1. In the following sections, we evaluate the usefulness of the upper bound in more realistic settings.
For dataset notation, we follow [8]. Each dataset is labeled as n_samples-n_features-centers-cluster_std, indicating the parameter values used to generate the points. For example, 400-16-3-4 denotes a dataset of $400$ $16$-dimensional vectors partitioned into $3$ clusters with standard deviation $4$, which controls within-cluster dispersion. For these datasets, we use Euclidean distance.
Table 2: Experimental results on synthetic data. Comparison of clustering performance and the proposed upper bound. Avg. distance indicates the average pairwise distance between cluster centers.
Dataset $K$ ARI AMI $\operatorname{ASW}$ $\texttt{UB}(\Delta)$ $1-\frac{\operatorname{ASW}}{\texttt{UB}(\Delta)}$ Avg. distance
400-64-5-6 5 1.0 1.0 $0.249$ $0.376$ $0.339$ $65.2$
400-64-2-2 2 1.0 1.0 $0.673$ $0.673$ $0.000$ $64.5$
400-128-7-3 7 1.0 1.0 $0.522$ $0.566$ $0.078$ $92.3$
1000-300-5-2 5 1.0 1.0 $0.668$ $0.678$ $0.015$ $143.7$
10000-32-20-2 20 1.0 1.0 $0.626$ $0.774$ $0.192$ $47.0$
10000-1024-20-4 20 1.0 1.0 $0.417$ $0.454$ $0.081$ $263.1$
In Table 2, we apply the upper bound to six synthetic datasets and compare the results with PAMSIL clustering. For dataset 400-64-2-2, the upper bound at $0.673$ confirms that PAMSIL reaches a globally optimal silhouette score. For three of the remaining datasets, $\texttt{UB}(\Delta)$ proves that the PAMSIL $\operatorname{ASW}$ is within 8 $\%$ of the global maximum. For dataset 400-64-5-6 with $\texttt{UB}(\Delta)=0.376$, we record an empirical $\operatorname{ASW}$ of $0.249$. In this case, it remains unclear whether further substantial improvements of the $\operatorname{ASW}$ are possible. However, interpreting the attained value in the interval [1,0.376][-1,0.376] enables more well-founded conclusions compared to using the standard range [1,1][-1,1]. Lastly, the reported ARI and AMI scores tell us that PAMSIL finds the ground truth clustering in all six datasets.
Figure 3 illustrates why it is necessary to scan all values of $k$ when constructing the upper bound in Proposition 1. In this example, we have picked a random data point $i$ from the dataset 1000-300-5-2, and the graph shows the quantity $1-q(i,\Delta,k)$ over a range of $k$-values, demonstrating that, for example, $k=2$ does not correspond to a valid bound in this case. Instead, the minimum of $q(i,\Delta,k)$ occurs at $k=200$, which coincides with the generated cluster sizes. Although in real-world datasets with less clearly separated clusters, we generally expect the difference between the average distance to the n2n-2 farthest neighbors and the nearest-neighbor distance to be sufficiently large that $k=2$ minimizes $q$. We suspect that this behavior often holds true even for data points whose cluster size in $\mathcal{C}^{*}$ is much larger than 2. Assuming that this speculation is true, it would be very unlikely that $\texttt{UB}(\Delta)$ is actually close to sharp in real-world settings.
Figure 3: For a random data point in the synthetic dataset 1000-300-5-2, we show the quantity $1-q(\Delta,k)$ for a range of $k$-values. The red dashed horizontal line shows the point’s silhouette width achieved by PAMSIL clustering with $K=5$. The maximum of $1-q(\Delta,k)$ is the upper bound from Proposition 1. The maximum is given at $k=200$, which corresponds to the cluster sizes in this dataset.
Figure 4 shows runtime measurements for Algorithm 1 (mm=1). Additionally, we include the runtime of constructing $\hat{\Delta}$ from $\Delta$, i.e., the second row in Algorithm 1. As a reference, we also report the time needed to compute the dissimilarity matrix $\Delta$ (although $\Delta$ is assumed to be precomputed in Algorithm 1). The experiments were conducted on an Apple MacBook Pro (M1, 2020) with 16GB of RAM, and the software environment was Python 3.10. Distance matrices of varying size and float32 data type were generated through the NumPy and scikit-learn packages. Specifically, we used the random.randn function to generate $n$ samples with $32$ features (seed fixed to 0), and the metrics.pairwise_\_distances function to generate all pairwise distances. The runtime of Algorithm 1 does not depend on the degree of clustering structure present in the dataset, which is why make_\_blobs was not used for this task. For each distance matrix, the programs were executed five times. The graph shows the average runtimes for each number of samples $n$, and the error bars indicate one standard deviation. It is clear from the figure that in our implementation, a substantial amount of the runtime of Algorithm 1 is spent on sorting each of the $n$ rows of $\Delta$ to obtain $\hat{\Delta}$.
Figure 4: Runtime experiments on synthetic data (log-linear plot).
5.2 UCI datasets
We evaluate our method on eight real-world datasets from the UCI Machine Learning Repository [13], commonly used in clustering and classification research. Table 3 summarizes their key characteristics along with the distance metrics employed. Standard preprocessing steps, including feature scaling and outlier removal, are applied where appropriate. Each dataset provides a known number of classes, which we use to determine the number of clusters $K$. When all class labels are available, we report the ARI and AMI to verify that these datasets represent meaningful clustering structures.